OpenStudy (unofficialllyy):

Your car has stalled near a gas station. You and your friend get out and attempt to push it. You push with a force of 120N, and your friend pushes in the same direction as you with a force of 100N. Friction also acts on the car with a force of 20N. If the car has a mass of 1,000kg, what is its acceleration?

OpenStudy (unofficialllyy):

@sweetburger

OpenStudy (brainlesslonelygirl):

you and your friend pushing, total force = 120+100=220 frictional force in opposite direction = 220-20= 200 equation for force, F= ma mass is calculated in grams hence we need to convert = 1000kg to g is 1000*1000=1000000 g now making acceleration the subject, a= F/m a=200/1000000, a= 2*10^-4 m/s^2